Draw the normal, shear and moment diagrams for the following problems 10 kN 10 kN 10 KN F-1 m – 1 m –1 m-- 2 m F-1 m-
Answer #1

Solo 10kN lokn lokN Ax KSK nim im Tim Tim 2m -- -- i © XV IM I Cy Efyzo, На + (= 1o EN EMA 2-0, Gyxu = 10X1+10x2 » Cy= 15kn|Jokn JOIN TOWN I MB Ay im im 0=15kn (7.1 Y-SEN Lo $© *cy=15kn from FBD-1 Efyzo, Ay+ By = 30 kN EMB2=0, Ayx3- 10x2- 10 x1 = MBNormal force diagram :- 13.416 KN Shear force diagram :- 19 6700 8t)Bending moment diagram : 15_ SISEN-M 15 KN-M - a